You have 10 posters.

You know 10 people such that 1 person is willing to pay $9 for the poster, another $8, another $7, each successive person is willing to pay one dollar less all the way down to a 10th person willing to take the poster for free.

If you can only charge one price, what price do you charge?
Answer #1

Answer

You will charge a price which will maximize revenue as the costs are not provided.

A firm's goal is to maximize profit but if there are no costs then it is revenue maximization.

total revenue =P*Q

at a price of $9 only 1 person is going to buy it and so on to at zero price all can buy.

TR(10)=10*0=0, TR(9)=1*9=9 and so on,

Q P TR
10 0 0
9 1 9
8 2 16
7 3 21
6 4 24
5 5 25
4 6 24
3 7 21
2 8 16
1 9 9

the total revenue is maximum at Q=5 units you will charge $5 if you can charge only one price

P=$5